OpenStudy (anonymous):

Determine if the lines 9x-2y=16 and 2x-9y= -18 are parallel, perpendicular, or neither……I got neither

OpenStudy (anonymous):

if the lines are parallel. their slopes must be equal. if they are perpendicular their slopes product will be -1 can you find the slopes of above lines ?

OpenStudy (anonymous):

for the slopes i got -9 and -2

OpenStudy (anonymous):

oops !! slopes you found are not correct. for the line in the form ax+by=c slope is m=-a/b for example if i have 3x+2y=c then here a=3 ,b=2 slope will be m=-(3)/2= -3/2 can you find the slopes now ?

OpenStudy (anonymous):

soo the answer wouldnt be neither??

OpenStudy (anonymous):

can you find the slopes ? find them let me know.